# Evaluating Bounded Superintelligent Authority in Multi-Level Governance: A Framework for Governance Under Radical Capability Asymmetry

### 03 Apr 2026 v1

@@ 6.5 Subsidiarity
The allocative dimension of subsidiarity [38, 54] does not oppose a higher-tier AI governance role; it supports one, for problems that genuinely exceed lower-tier capacity. The structural tension lies in scope limitation, and it operates through at least three mechanisms. The first is voluntary legislative expansion: lower-tier legislatures, observing the higher tier’s superior performance in its initial domain, rationally delegate additional domains through ordinary constitutional processes. Each individual delegation is democratically authorized, but the cumulative effect is the same erosion of lower-tier authority. The second is unauthorized expansion via crisis response: an acute crisis (pandemic, financial collapse, military threat) creates pressure for the higher tier to act beyond its mandate, and the superior capability that makes its intervention effective also makes post-hoc ratification likely. The third is inherently fuzzy domain boundaries: complex governance problems resist clean jurisdictional division, and an agent capable of identifying interdependencies across domains has a structural incentive to expand its purview in the name of coordination.
@@ 6.7 Synthesis: structural versus contingent failures
Having examined each dimension individually, the analysis now classifies the failure modes across all six. Three categories emerge. Contingent failures depend on current technical limitations and could in principle be resolved without altering the case’s core features. These include interpretability (an active research area), corrigibility (whose classification depends on progress in architectural approaches that remain undemonstrated at the relevant capability level), and specific accountability mechanisms (designable given sufficient interpretability). A caveat is needed for accountability. While the individual components are contingent on technical progress, the correlated degradation of all three components under radical asymmetry, the qualitative shift from independent to linked failures, has a structural character that the “contingent” label may understate.
Among the structural failures, which persist across implementations and intensify as the capability gap widens, a further distinction is useful. Some structural failures are design-tractable: they arise from specific architectural choices and might yield to institutional innovation without requiring new normative theory. The single-point-of-failure problem and the subsidiarity scope-limitation problem fall into this category. A multi-agent architecture with mutual checks could address the first; justiciable mandate boundaries enforced by independent review could address the second. These are hard institutional design problems, but they are problems of the kind governance theory already knows how to think about.
@@
In existing governance institutions, the six dimensions function largely as independent checks: an institution can fail on accountability while satisfying non-domination requirements, or strain subsidiarity while maintaining corrigibility. The analysis in Section 4 demonstrates this independence across multiple institutional examples. Under radical capability asymmetry, this independence breaks down. Legitimacy requires that oversight bodies comprehend the agent’s reasoning (public reason), accountability requires that they evaluate its performance (answerability), and non-domination requires that they contest its decisions (contestatory control). All three presuppose the cognitive comparability between oversight body and agent: when that comparability fails, all three degrade together. The transition from bounded to radical asymmetry is thus not merely a quantitative increase in difficulty along each dimension; it is a qualitative shift from independent to correlated failure modes.
@@
A final observation connects the analysis to the case’s own exclusions. Section 5 specified that the prospective case is not world government, not unrestricted authority, and not a replacement for democratic self-governance. The analysis suggests that maintaining these exclusions is the central difficulty. The subsidiarity analysis shows how scope limitation erodes under pressure from demonstrated superior capability. The corrigibility analysis shows how constitutional bounds require enforcement against an agent capable of circumventing them. The non-domination analysis shows how formal democratic authorization becomes nominal when the governed cannot meaningfully contest the governor’s conduct. The prospective case describes a bounded arrangement; the analysis suggests that the boundaries are load-bearing in a way that the capability asymmetry systematically undermines.
@@ 7 Discussion
This paper contributes to political theory, not to predictions about ASI development. The evaluation framework is designed for broad use; the extreme prospective case makes each dimension’s demands visible and exposes assumptions that milder cases leave untested. The framework applies to any proposal placing an artificial agent in a governance role, including algorithmic regulatory systems [90, 28], AI-assisted policy optimization, and automated adjudication systems.
@@
A circularity objection must be addressed directly. The framework’s dimensions are drawn from democratic and republican political theory, which embeds commitments to public reason, contestatory control, and cognitive accessibility. Applying such a framework to an entity defined as cognitively exceeding all humans risks producing a negative result by construction. Two responses address this. First, the framework produces varied and non-predetermined results when applied to existing institutions in Section 4: algorithmic sentencing partially fails on legitimacy, central banks largely satisfy accountability, emergency powers strain corrigibility. A rigged framework would fail these cases uniformly; instead, it discriminates among arrangements in ways that track real analytical distinctions. The negative result for the ASI case is earned by the analysis, not built into the premises. Second, rather than denying the framework’s commitments, the productive response is to make them explicit and ask, for each failure, whether it reveals a problem with the prospective case or a limitation of the framework itself. Where the prospective case fails on public reason grounds, this exposes governance theory’s dependence on cognitive comparability, an assumption that has gone unexamined because it has always been satisfied. The negative assessment and the theoretical insight are linked: the prospective case works as a stress test precisely because it violates assumptions the theory normally takes for granted.
@@
Previous extensions of Rawlsian and republican frameworks to new domains (international justice, intergenerational obligations, the moral status of non-human animals [21]) have all preserved the assumption of cognitive comparability intact. The ASI case is the first that breaks it. These frameworks were not built for a world where the governed cannot comprehend the reasoning of those who govern them. No amount of careful application bridges a gap the theory’s foundations do not address. The analysis opens a foundational question for governance theory.
@@
The failure taxonomy (Table 2) identifies two problems that need new theory, not just better institutional design. The first is the public reason problem under cognitive incomprehensibility: what does legitimate justification look like when the justifying agent’s reasoning exceeds the cognitive reach of those it is addressed to? The second is the non-domination problem under permanent capability asymmetry: what does contestatory control mean when the agent being contested radically outperforms the contesting parties? These are not questions existing governance theory answers badly; they are questions it has not needed to ask. They will become pressing well before the superintelligence case, as increasingly capable AI systems take on advisory and decision-support roles that strain the limits of human oversight.
Power concentration and prior work
Any analysis of ASI governance might be read as legitimating the development trajectory. The present analysis supports the opposite conclusion. Four of six dimensions show that bounded superintelligent authority cannot meet the conditions democratic and republican theory require. The framework does not assume ASI will or should be developed; it asks what governance theory reveals when the prospective case is taken seriously.
Bostrom et al. [11] identify policy desiderata for superintelligent AI through a vector field approach. This paper extends that approach by specifying concrete dimensions from governance theory, setting necessary conditions that allow pass/fail evaluation rather than directional guidance, and sorting failures into categories that distinguish solvable design problems from those requiring new theory.
@@ 8 Conclusion
The analysis suggests several directions for further work. The structural failures identified here could be reframed as design constraints, bounding the space of viable proposals rather than ruling out bounded AI governance entirely. The non-domination and single-point-of-failure problems suggest that any viable proposal would need a multi-agent architecture with mutual checks, while the public reason problem suggests the agent’s domain must be restricted to areas where outcome-based assessment is feasible. Formal models of delegation under radical capability asymmetry could test whether institutional design can maintain the monitoring capacity that delegation requires, even when no individual can comprehend the agent [57]. Work connecting governance theory with constitutional AI research, where publicly deliberated constitutions have been shown to guide AI behavior as effectively as expert-designed ones [8, 52], could identify ways to preserve contestatory control when direct cognitive oversight is not possible. The goal is to determine whether these structural failures are hard limits or problems that yield to creative institutional design.
A framework that could reach more permissive conclusions would need to ground legitimacy in something other than public reason (perhaps pure outcome assessment), ground non-domination in something other than contestatory control (perhaps structural constraints sufficient to prevent arbitrary interference regardless of the agent’s capability), and show that the correlated failure pattern identified here can be broken apart through institutional design. Whether such a framework can be built without abandoning commitments central to democratic governance is itself worth investigating.
Governance theory has always developed in response to challenges that initially appeared to exceed the capacity of existing institutional forms. The challenge posed by radical capability asymmetry is distinctive not because it is unprecedented in kind (delegation under information asymmetry is as old as governance itself) but because it forces the discipline to confront assumptions it has never needed to state. The analysis identifies which of those assumptions are load-bearing and why they break down: the capability asymmetry undermines the monitoring, evaluation, and sanctioning capacities on which delegated governance depends. Whether these assumptions are necessary commitments of governance theory or contingent features of a world in which all governance agents have been human is a question that political theorists, legal scholars, AI researchers, and institutional designers are better positioned to answer together than any single discipline alone. The framework and failure taxonomy offered here are intended as tools for that work.

### 14 Apr 2026 v2

@@ 6.5 Subsidiarity
The allocative dimension of subsidiarity [38, 54] does not oppose a higher-tier AI governance role; it supports one, for problems that genuinely exceed lower-tier capacity. The structural tension lies in scope limitation, and it operates through at least three mechanisms. The first is voluntary legislative expansion: lower-tier legislatures, observing the higher tier’s superior performance in its initial domain, rationally delegate additional domains through ordinary constitutional processes. Each individual delegation is democratically authorized, but the cumulative effect is the same erosion of lower-tier authority.
The second is unauthorized expansion via crisis response: an acute crisis (pandemic, financial collapse, military threat) creates pressure for the higher tier to act beyond its mandate, and the superior capability that makes its intervention effective also makes post-hoc ratification likely. The third is inherently fuzzy domain boundaries: complex governance problems resist clean jurisdictional division, and an agent capable of identifying interdependencies across domains has a structural incentive to expand its purview in the name of coordination.
@@ 6.7 Synthesis: structural versus contingent failures
The failure modes across all six dimensions (Figure 1) fall into three categories. Contingent failures depend on current technical limitations and could in principle be resolved without altering the case’s core features. These include interpretability (an active research area), corrigibility (whose classification depends on progress in architectural approaches that remain undemonstrated at the relevant capability level), and specific accountability mechanisms (designable given sufficient interpretability). A caveat applies to accountability: while the individual components are contingent on technical progress, the correlated degradation of all three under radical asymmetry has a structural character that the “contingent” label may understate.
Among the structural failures, which persist across implementations and intensify as the capability gap widens, a further distinction is useful. Some structural failures are design-tractable, arising from specific architectural choices that might yield to institutional innovation without new normative theory. The single-point-of-failure problem and the subsidiarity scope-limitation problem fall into this category. A multi-agent architecture with mutual checks could address the first; justiciable mandate boundaries enforced by independent review could address the second. These are hard institutional design problems, but they are problems of the kind governance theory already knows how to think about.
@@
In existing governance institutions, the six dimensions function largely as independent checks. An institution can fail on accountability while satisfying non-domination requirements, or strain subsidiarity while maintaining corrigibility, and Section 4 shows this independence across several examples. Under radical capability asymmetry, the independence breaks down. Three dimensions quietly share a common prerequisite. Legitimacy on the public-reason view depends on the oversight body’s being able to comprehend the agent’s reasoning. Accountability, through its answerability component, depends on that body’s being able to evaluate the agent’s performance. Non-domination in Pettit’s sense depends on its being able to contest the agent’s decisions. Each of those capacities is a special case of the same underlying condition, and when cognitive comparability fails, all three degrade together. The transition from bounded to radical asymmetry is not really a matter of each dimension becoming harder in degree; it is the loss of a background condition that used to keep the dimensions independent.
@@
These results bear directly on the case’s own exclusions. Section 5 specified that the prospective case is not world government, not unrestricted authority, and not a replacement for democratic self-governance. Maintaining these exclusions is the central difficulty. Scope limitation erodes under pressure from demonstrated superior capability. Constitutional bounds require enforcement against an agent capable of circumventing them. Formal democratic authorization becomes nominal when the governed cannot meaningfully contest the governor’s conduct. The prospective case describes a bounded arrangement, but the boundaries are load-bearing in a way that the capability asymmetry systematically undermines.
@@ 7 Discussion
This paper contributes to political theory, not to predictions about ASI development. The framework does not presuppose that ASI will or should be developed; it asks what governance theory reveals when the prospective case is taken seriously. The extreme case makes each dimension’s demands visible and exposes assumptions that milder cases leave untested, but the dimensions apply to any proposal placing an artificial agent in a governance role, including algorithmic regulatory systems [90, 28], AI-assisted policy optimization, and automated adjudication systems. The opacity/incomprehensibility distinction already applies to regulatory oversight of large language models, whose outputs can be inspected but whose internal reasoning cannot be fully traced. The correlated failure pattern, where independent accountability checks degrade together, applies wherever the capability gap between a governed system and its overseers widens faster than interpretive tools can close it. These are not future problems.
@@
A circularity objection must be addressed directly. The framework’s dimensions are drawn from democratic and republican political theory, which embeds commitments to public reason, contestatory control, and cognitive accessibility. Applying such a framework to an entity defined as cognitively exceeding all humans risks producing a negative result by construction. Two responses address this. First, the framework produces varied and non-predetermined results when applied to existing institutions in Section 4: algorithmic sentencing partially fails on legitimacy, central banks largely satisfy accountability, emergency powers strain corrigibility. A rigged framework would fail these cases uniformly; this one discriminates among arrangements in analytically meaningful ways. The negative result for the ASI case is earned by the analysis, not built into the premises. Second, rather than denying the framework’s commitments, the productive response is to make them explicit and ask, for each failure, whether it reveals a problem with the prospective case or a limitation of the framework itself. Where the prospective case fails on public reason grounds, this exposes governance theory’s dependence on cognitive comparability, an assumption that has gone unexamined because it has always been satisfied. The negative assessment and the theoretical insight are linked: the prospective case works as a stress test precisely because it violates assumptions the theory normally takes for granted.
@@
Previous extensions of Rawlsian and republican frameworks to new domains (international justice, intergenerational obligations, the moral status of non-human animals [21]) have all preserved the assumption of cognitive comparability. The ASI case is the first that breaks it. These frameworks were not built for a world where the governed cannot comprehend the reasoning of those who govern them, and no amount of careful application bridges a gap the theory’s foundations do not address.
@@
The failure taxonomy (Table 2) identifies two problems that need new theory rather than better institutional design. One is the public reason problem under cognitive incomprehensibility. What does legitimate justification even look like once the justifying agent’s reasoning exceeds the cognitive reach of the people it is addressed to? The other is the non-domination problem under permanent capability asymmetry. What is contestatory control supposed to mean when the agent being contested radically outperforms the contesting parties? Existing governance theory does not answer these badly. It has just not had reason to ask them before. Both will become pressing well before the superintelligence case itself arrives, as increasingly capable AI systems take on advisory and decision-support roles that stretch the limits of human oversight.
@@ 8 Conclusion
The structural failures identified here open several directions for further work. They could be reframed as design constraints, bounding the space of viable proposals rather than ruling out bounded AI governance entirely. The non-domination and single-point-of-failure problems suggest that any viable proposal would need a multi-agent architecture with mutual checks, while the public reason problem suggests the agent’s domain must be restricted to areas where outcome-based assessment is feasible. Formal models of delegation under radical capability asymmetry could test whether institutional design can maintain the monitoring capacity that delegation requires, even when no individual can comprehend the agent [57]. Work connecting governance theory with constitutional AI research, where publicly deliberated constitutions have been shown to guide AI behavior as effectively as expert-designed ones [8, 52], could identify ways to preserve contestatory control when direct cognitive oversight is not possible. The goal is to determine whether these structural failures are hard limits or problems that yield to creative institutional design.
A framework that could reach more permissive conclusions would have substantial work to do. It would need a basis for legitimacy other than public reason, perhaps something closer to pure outcome assessment. It would need a basis for non-domination other than contestatory control, possibly structural constraints strong enough to prevent arbitrary interference regardless of the agent’s capability. And it would have to show that the correlated failure pattern identified above can be pulled apart through institutional design. Whether any such framework can be built without abandoning commitments central to democratic governance is itself an open question.
Governance theory has always developed in response to challenges that initially appeared to exceed the capacity of existing institutional forms. The challenge posed by radical capability asymmetry is distinctive not because it is unprecedented in kind (delegation under information asymmetry is as old as governance itself) but because it forces the discipline to confront assumptions it has never needed to state. The capability asymmetry undermines the monitoring, evaluation, and sanctioning capacities on which delegated governance depends, and the framework reveals which foundational assumptions are load-bearing and why they break down. Whether these assumptions are necessary commitments of governance theory or contingent features of a world in which all governance agents have been human is a question that political theorists, legal scholars, AI researchers, and institutional designers are better positioned to answer together than any single discipline alone. The framework and failure taxonomy offered here are intended as tools for that work.